Fill in the blanks with >,
 (–3) + (–6) ______ (–3) – (–6) 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(-3) + (-6) = -9
(-3)–(-6) = -3 + 6 =3
-9 < 3
Therefore (-3) + (-6) > (-3)–(-6)

Thanks